OpenStudy (1645323):

The weekly revenue for a company is R = -3p2 + 60p + 1060, where p is the price of the company’s product. Use the quadratic formula to find the price for which the weekly revenue would be $1000. Select one: a. $10.00 b. $15.25 c. $20.95 d. $18.90

OpenStudy (1645323):

@hass33 @le0n

OpenStudy (anonymous):

\[-3p^{2}+60p+1060=1000\]\[-3p^{2}+60p+60=0\]\[p^{2}-20p-20=0\]